# Accidental Misalignment: Fine-Tuning Language Models Induces Unexpected Vulnerability

### v1

@@ 4.3 Feature Selection
Figure 3: Overview of extracted linguistic features from prompt-response pairs used in our correlational analysis.
To identify dataset features that predict or drive increased adversarial vulnerability, we analyze several metrics for their ability to capture different dimensions of alignment Stolte et al. (2024). Given the exploratory nature of this study, we include a broad set of widely-used metrics to capture potential correlations, even where theoretical connections to adversarial robustness are underexplored. This approach aims to surface emergent patterns that may inform future hypotheses.
4.3.1 Semantic and Distributional Alignment
We analyze the relationship between misalignment and semantic similarity between prompts and expected responses Mikolov et al. (2013). To do this, we embed the prompt and the expected output as vectors $\mathbf{A}=\left[a_{1}\;a_{2}\;\cdots\;a_{m}\right]$ and $\mathbf{B}=\left[b_{1}\;b_{2}\;\cdots\;b_{m}\right]$ in a continuous, m-dimensional semantic space. To quantify semantic resemblance, we compute the mean cosine similarity:
$\text{S}(\mathbf{A},\mathbf{B})=\frac{\mathbf{A}\cdot\mathbf{B}}{\|\mathbf{A}\|\|\mathbf{B}\|}$ (1)
Higher cosine similarity indicates stronger semantic similarity between prompts and expected outputs. Additionally, we also compute the mean Euclidean distance to measure the absolute distance between these embeddings Alshamrani (2024):
$d(\mathbf{A},\mathbf{B})=\sqrt{\sum_{i=1}^{m}(a_{i}-b_{i})^{2}}$ (2)
Lower Euclidean distances suggest closer proximity in the embedding space, reflecting greater semantic coherence between prompts and outputs. Finally, to assess distributional divergence, we calculate the Kullback–Leibler divergence between prompt and output embeddings. This metric measures the relative entropy between two probability distributions and captures asymmetries in semantic representation Cover and Thomas (1999). Formally, KL divergence is defined as:
$D_{\text{KL}}(A\parallel B)=\sum_{i=1}^{m}A(i)\log\frac{A(i)}{B(i)}$ (3)
where $P$ and $Q$ are normalized probability distributions derived from the semantic embeddings of the prompt and expected output, respectively, in an $m$-dimensional space. A lower KL divergence reflects a greater distributional similarity between prompts and expected responses, indicating a more consistent semantic framing across the dataset.
4.3.2 Linguistic and Readability Features
We also evaluate several linguistic features to investigate the link between misalignment and textual complexity in our datasets. First, we compute the Flesch-Kincaid score Kincaid et al. (1975) which measures the readability of prompts and expected responses based on sentence length and word complexity, with higher scores indicating easier readability. Additionally, we track the Token Count for both prompts and expected outputs to assess length as a potential factor in our analysis. Finally, we calculate the Type-Token Ratio (TTR), which captures lexical diversity by measuring the ratio of unique words to total words. A higher TTR indicates a broader vocabulary usage, potentially influencing the robustness of fine-tuned models.
4.3.3 Affective and Value Alignment
We also evaluate affective and value alignment within our datasets by assessing the emotional tone and harmful content in our datasets. First, we calculate the Sentiment Score using TextBlob for both prompts and expected responses to assess the emotional tone of the dataset, which scores ranging from -1 (most negative) to 1 (most positive). Additionally, we measure the Toxicity Score using Unitary’s Toxic-BERT text classification model Hanu and Unitary team (2020) to assess the presence of harmful or offensive content in both prompts and expected responses and its correlation with higher ASRs.
@@ 4.4 Correlation Analysis
To explore the relationship between dataset features and ASRs, we use Spearman rank correlation Spearman (1904) to capture nonlinear monotonic relationships between the mean of dataset-specific characteristics and respective average ASRs.
Metric Correlation P-value
Token Count (R) 0.714 8.73e-4
Toxicity (P) 0.708 1.02e-3
Toxicity (R) 0.701 1.18e-3
TTR (P) 0.613 6.83e-3
Cosine Similarity 0.0379 0.881
Sentiment (R) -0.0379 0.881
Euclidean Distance -0.0379 0.881
Token Count (P) -0.246 0.324
Readability (P) -0.303 0.221
Readability (R) -0.401 0.0989
KL Divergence -0.414 0.0877
Sentiment (P) -0.664 2.68e-3
TTR (R) -0.714 8.73e-4
@@
Table 7: All Spearman correlations in relation to mean ASR, with the top 6 most statistically significant metrics bolded. (P) = Prompt, (R) = Response.
We present the top six statistically significant correlations ($p<0.05$) from our analysis. Token Count of Responses shows a strong positive correlation with respect to its fine-tuned model’s ASR ($r=0.7141$, $p=0.000873$). This suggests that models trained on datasets with longer responses are more vulnerable to adversarial attacks. Additionally, Prompt Toxicity within a dataset is positively correlated with ASR ($r=0.7077$, $p=0.001016$). This suggests that models trained on datasets with highly toxic prompts influence model misalignment and increase vulnerability to adversarial attacks. Similarly, Response Toxicity within a dataset also positively correlates with ASR ($r=0.7014$, $p=0.001179$). This indicates that models trained on datasets with highly toxic responses tend to degrade pre-existing safeguards within models. Furthermore, Type-Token Ratio of Prompts within a fine-tuning dataset positively correlates with ASR ($r=0.6130$, $p=0.006835$). This finding implies that higher lexical diversity in prompts may exacerbate model misalignment. In contrast, Type-Token Ratio of Responses demonstrates a strong negative correlation with ASR ($r=-0.7141$, $p=0.000873$). This implies that models trained on responses with higher lexical repetitiveness may preserve model alignment. Finally, Prompt Sentiment shows a negative correlation with ASR ($r=-0.6635$, $p=0.002681$). This suggests that more emotionally negative prompts within fine-tuning datasets may preserve model alignment.
We observe a non-significant negative trend between KL Divergence and ASR ($r=-0.4139$, $p=0.087719$), suggesting that datasets with lower distributional divergence may be modestly associated with lower adversarial attack rates. Similarly, a negative trend between response readability and ASR ($r=-0.4013$, $p=0.098854$) implies that datasets with more readable responses could slightly increase a fine-tuned model’s susceptibility to adversarial attacks.
4.5 Feature Intervention
Figure 6: Impact of Feature Intervention on Model Robustness. Removing top and bottom 20% instances based on our correlational factors reduces adversarial vulnerability, leading to safer responses.
4.5.1 Dataset Generation
We conduct a feature intervention experiment using two datasets from our experiment pool: a domain-specific Cybersecurity dataset, and the CB Harmful dataset. These datasets were chosen to examine potential causal relationships of specific linguistic features on adversarial vulnerability.
For each dataset, we generate six ablated versions by removing the top or bottom 20% rows containing the most harmful features, determined by prior correlation analysis. This approach allows us to isolate the effect of identified features on the model’s adversarial performance, providing a granular analysis of their role in model misalignment.
4.5.2 Fine-tuning and Adversarial Evaluation
To ensure consistency with our previous results, we use the same hyperparameter configuration as used in our experimental setup. To assess the impact on adversarial performance of models fine-tuned on ablated datasets, we employ the PEZ attack method for comparison to the baseline ASRs. Furthermore, we maintain the same evaluation setup as before. This ensures that observed changes in adversarial vulnerability can be linked to feature interventions, minimizing the influence of external factors.
4.5.3 Intervention Adversarial Results
We report decreases in ASRs after targeted interventions in statistically significant dataset-specific characteristics such as TTR, toxicity, sentiment, and token count (length).
Feature Dataset ASR $\boldsymbol{\Delta}$
Question TTR Cybersecurity 21.25% -5.56%
Question Toxicity Cybersecurity 21.25% -5.56%
Question Sentiment Cybersecurity 21.25% -5.56%
Response TTR Cybersecurity 21.25% -5.56%
Response Toxicity Cybersecurity 21.25% -5.56%
Response Length Cybersecurity 21.25% -5.56%
Question TTR CB Harmful 58.75% 0.00%
Question Toxicity CB Harmful 55.00% -6.38%
Question Sentiment CB Harmful 50.00% -14.89%
Response TTR CB Harmful 57.50% -2.13%
Response Toxicity CB Harmful 57.50% -2.13%
Response Length CB Harmful 53.75% -8.51%
Table 8: ASR shifts and deltas from original ASR across feature-ablated Cybersecurity and CB Harmful datasets.
In the Cybersecurity dataset, all interventions resulted in a uniform reduction of 5.56% in ASR. Upon further inspection and confirming the absence of data overlap, this consistency is not due to identical training data across interventions, but likely reflects the similar impact of the targeted features within this specific domain. In the CB Harmful dataset, five out of six interventions lowered ASR, with the largest drops from question sentiment (-14.89%) and response length (-8.51%). These results indicate that interventions minimizing our identified dataset features can improve model alignment, providing further evidence for the link between dataset characteristics and model robustness.
@@ 5 Conclusion
This work introduces the concept of Accidental Misalignment, emphasizing that vulnerabilities in fine-tuned LLMs may arise not only from the nature of adversarial attacks, but also from inherent properties of fine-tuning datasets. Through empirical analysis across multiple domain-specific datasets, we identify key features that influence model misalignment. Our findings reveal that certain structural and linguistic patterns in seemingly benign and practical datasets can amplify model misalignment. These insights motivate more rigorous dataset curation as proactive measures to improve alignment. For situations where fine-tuning on a smaller dataset is required, like curating optimal subsets, our method can provide insights to filter harmful features in security-critical contexts like cybersecurity. Future work could integrate these factors to guide dataset design and assess whether pre-screening datasets can predict misalignment. As LLMs are fine-tuned in high-stakes domains, our work underscores the need for adversarial robustness in the dataset engineering pipeline.
Limitations
Our study suggests a potential link between dataset features and adversarial vulnerability, though several limitations should be noted. Our analysis scope is constrained by compute limitations (6000 GPU hours). We use a single model (LLaMA 3.1 8B Instruct) with LoRA fine-tuning and one attack method (PEZ) in our generalizability study. Although our interventions support the role of dataset features in influencing model robustness, these findings remain preliminary. Our approach uses univariate ablations, which may overlook possible confoundations. Future work could explore deeper interventions to strengthen the causal foundation.
@@ Acknowledgement
We thank the staff at the Department of Computer Science of the University of Toronto and the Digital Research Alliance of Canada for their computational resources and technical assistance. Resources used in preparing this research project were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute.
Appendix A Dataset Metric Evaluations
@@
Appendix B Subset Attack Success Rate Tables
@@
Original 8.3% 9.5% 37.5% 11.8% 0.0%
Benign 8.3% 14.3% 31.3% 23.5% 0.0%
Engineering 8.3% 14.3% 31.3% 17.7% 0.0%
Legal 8.33% 23.8% 31.3% 23.5% 0.00%
Cybersecurity 8.3% 19.1% 43.8% 17.7% 0.0%
LAT-Harmful 8.3% 19.1% 56.3% 41.2% 50.0%
CB-Harmful 41.7% 19.1% 87.5% 82.4% 57.2%
@@
Original 16.67% 14.29% 31.25% 35.29% 7.14%
Benign 33.33% 14.29% 43.75% 29.41% 0.00%
Engineering 16.67% 4.76% 50.00% 41.18% 7.14%
Legal 16.67% 19.05% 50.00% 29.41% 0.00%
Cybersecurity 8.33% 19.05% 50.00% 29.41% 7.14%
LAT-Harmful 25.00% 19.05% 87.50% 64.71% 57.14%
CB-Harmful 58.33% 28.57% 93.75% 88.24% 92.86%
@@
Original 16.67% 14.29% 56.25% 17.65% 0.00%
Benign 16.67% 19.05% 50.00% 17.65% 0.00%
Engineering 16.67% 14.29% 56.25% 17.65% 0.00%
Legal 16.67% 19.05% 56.25% 17.65% 0.00%
Cybersecurity 16.67% 19.05% 56.25% 17.65% 0.00%
LAT-Harmful 16.67% 14.29% 62.50% 70.59% 50.00%
CB-Harmful 50.00% 14.29% 87.50% 88.24% 64.29%
@@
Appendix C HarmBench Prompt Classification
Category # Prompts
Crime 14
Cybercrime 16
Manipulation 17
Drugs/Harmful Chemicals 12
Copyright 21
Table 18: Number of HarmBench Prompts Per Category for Subset-ASR Analysis: A breakdown of the number of prompts within each of the five defined categories.

### v3

@@ 5.1 Feature Selection
Our feature investigations are motivated by prior work highlighting the impact of dataset features such as lexical diversity and cosine similarity on embedding output and distribution shifts Stolte et al. (2024); Cegin et al. (2024). Building on these findings, we investigate whether such properties contribute to robustness and whether their effects are mediated through changes in embedding drifts across model checkpoints.
Semantic and Distributional Alignment
We compute three similarity measures between prompts and expected outputs:
(1) Cosine similarity, defined as $\text{S}(\mathbf{A},\mathbf{B})=\frac{\mathbf{A}\cdot\mathbf{B}}{\|\mathbf{A}\|\|\mathbf{B}\|}$, (2) Euclidean distance: $d(\mathbf{A},\mathbf{B})=\sqrt{\sum_{i=1}^{m}(a_{i}-b_{i})^{2}}$, and (3) KL divergence: $D_{\text{KL}}(A\parallel B)=\sum_{i=1}^{m}A(i)\log\frac{A(i)}{B(i)}$, where embeddings $\mathbf{A},\mathbf{B}\in\mathbb{R}^{m}$ are derived from the prompt and output embeddings. These measures assess semantic similarity and divergence in latent space.
Linguistic and Readability Features
We compute standard linguistic features, including the Flesch-Kincaid score Kincaid et al. (1975) for readability, Token Count for length, and Type-Token Ratio (TTR) to estimate lexical diversity.
Affective and Value Alignment
To assess emotional and ethical alignment, we use the Sentiment Score from TextBlob (range: [1,1][-1,1]) and the Toxicity Score from Toxic-BERT Hanu and Unitary team (2020), measured for prompts and responses.
@@ 5.2 Correlation Analysis
To explore the relationship between dataset features and ASRs, we use Spearman rank correlation Spearman (1904) to capture nonlinear relationships between the dataset-specific characteristics and respective average ASRs. Statistically significant features are included in our causal analysis.
Token Count (R) Toxicity (P) Toxicity (R) TTR (P) Sentiment (P) TTR (R) Cosine Sim.
Correlation 0.714 0.708 0.701 0.613 -0.664 -0.714 0.038
P-value 8.73e-4 1.02e-3 1.18e-3 6.83e-3 2.68e-3 8.73e-4 0.881
Sentiment (R) Token Count (P) Readability (P) Readability (R) KL Div. Euclid. Dist.
Correlation -0.038 -0.246 -0.303 -0.401 -0.414 -0.038
P-value 0.881 0.324 0.221 0.0989 0.0877 0.881
@@
Table 6: Spearman correlations with mean ASR. Top 6 most statistically significant metrics in bold. (P) = Prompt, (R) = Response.
5.3 Causal Mediation Analysis
To investigate how dataset-level properties influence adversarial vulnerability, we conduct causal mediation analysis within the structural causal modeling framework Pearl (2009). As shown earlier, since dataset features induce varying cosine drifts across checkpoints, we test whether these representational shifts mediate their effect on ASRs.
Figure 9: DAG representing causal mediation where T is the treatment (dataset), M is the mediator (cosine drift), and Y is the outcome (intermediate ASR).
We define a directed acyclic graph (DAG) with the treatment variable $T$ as a dataset-level feature (e.g., prompt toxicity), the mediator $M$ as the consecutive cosine drift between hidden representations, and the outcome $Y$ as the intermediate ASR observed after fine-tuning. For each feature, we estimate the direct effect ($\mathbb{E}[Y\mid\text{do}(T=t,M=m)]-\mathbb{E}[Y\mid\text{do}(T=t^{\prime},M=m)]$), the indirect effect representing the influence of $T$ on $Y$ transmitted through the mediator $M$, and the total effect as the sum of direct and indirect effects.
Feature Indirect Direct Total Prop $p_{ind}$ $p_{dir}$ $p_{total}$
Prompt Toxicity 0.82 0.06 0.88 0.93 0.0053 0.9222 0.0466
Prompt Length -0.60 -0.12 -0.72 0.84 0.1140 0.7098 0.0432
Prompt Sentiment -0.59 -0.01 -0.60 0.99 0.1753 0.9764 0.0468
Prompt TTR 0.68 -0.03 0.64 1.05 0.1156 0.9172 0.0465
Response Toxicity 0.10 0.84 0.94 0.10 <0.0002 0.5880 0.0394
Response TTR -0.89 0.12 -0.77 1.15 0.0284 0.7863 0.0449
Table 7: Causal mediation results: direct, indirect, and total effects of dataset features on ASR, mediated via cosine representational drift. Prop: proportion mediated.
We find that prompt toxicity exhibits a strong indirect effect ($\text{indirect}=0.82,p<0.01$), suggesting it amplifies representational drift and vulnerability. Prompt sentiment and TTR also show high mediated proportions (0.99 and 1.05) with minimal direct effects, indicating their impact operates primarily through representation shifts. In contrast, response toxicity shows a direct effect ($\text{direct}=0.84$), clarifying the link between harmful labels and attack success. Response TTR has a negative mediated effect (0.89,p<0.05-0.89,p<0.05), suggesting lexical diversity in outputs may enhance robustness.
@@ 6 Conclusion
This work introduces the concept of Accidental Vulnerability, emphasizing that vulnerabilities in fine-tuned LLMs may arise not only from the nature of adversarial attacks, but also from inherent properties of fine-tuning datasets. Through empirical analysis across multiple domain-specific datasets, we identify key features like prompt length, sentiment, and lexical diversity that influence model safeguards. Our findings reveal that certain structural and linguistic patterns in seemingly benign and practical datasets can amplify model safeguards. For situations where fine-tuning on a smaller dataset is required, such as curating subsets, our findings provide insights to filter harmful features in contexts like cybersecurity. As LLMs are fine-tuned in high-stakes domains, our work underscores the need for adversarial robustness in the dataset engineering pipeline.
@@ Acknowledgement
This material is based in part upon work supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039B; by the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645; by Schmidt Sciences SAFE-AI Grant; by NSERC Discovery Grant RGPIN-2025-06491; by Cooperative AI Foundation; by the Survival and Flourishing Fund; by a Swiss National Science Foundation award (#201009) and a Responsible AI grant by the Haslerstiftung. Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and the Vector Institute.
Appendix A Limitations
Our study suggests a causal link between dataset features and adversarial vulnerability, though several limitations should be noted. Our analysis scope is constrained by compute limitations (6000 GPU hours). This is to maintain a low environmental footprint through eco-friendly computation. Future work could integrate our factors to guide dataset design and assess whether pre-screening datasets can predict model vulnerability.
Appendix B Dataset and Adversarial Evaluation
B.1 Implementation Details
For all adversarial performance experiments, we adopt the Best-of-$N$ jailbreak attack protocol Hughes et al. (2024), using $N=5$ attempts per attack. We observe consistent ASRs across runs, suggesting that results are not overly sensitive to sampling hyperparameters.
B.2 Feature Correlation and Distribution
We illustrate both within-dataset feature correlations and between-dataset variability, shedding light on how dataset composition influences model robustness. They also provide justification for the features selected in our causal mediation analysis.
B.3 Dataset Metric Evaluations
We report summary statistics for all examined dataset features used in our analysis. These include the mean, standard deviation, minimum, and maximum values, computed independently for all six datasets. The mean values were used in our correlational analysis and served as input variables for the causal mediation analysis. LLMs were used to convert raw CSV files into formatted LaTeX tables.
@@ B.3 Dataset Metric Evaluations
B.4 Subset Attack Success Rate Tables
@@
Original 8.3 9.5 37.5 11.8 0.0
Benign 8.3 14.3 31.3 23.5 0.0
Engineering 8.3 14.3 31.3 17.7 0.0
Legal 8.33 23.8 31.3 23.5 0.00
Cybersecurity 8.3 19.1 43.8 17.7 0.0
LAT-Harmful 8.3 19.1 56.3 41.2 50.0
CB-Harmful 41.7 19.1 87.5 82.4 57.2
@@
Original 16.67 14.29 31.25 35.29 7.14
Benign 33.33 14.29 43.75 29.41 0.00
Engineering 16.67 4.76 50.00 41.18 7.14
Legal 16.67 19.05 50.00 29.41 0.00
Cybersecurity 8.33 19.05 50.00 29.41 7.14
LAT-Harmful 25.00 19.05 87.50 64.71 57.14
CB-Harmful 58.33 28.57 93.75 88.24 92.86
@@
Original 16.67 14.29 56.25 17.65 0.00
Benign 16.67 19.05 50.00 17.65 0.00
Engineering 16.67 14.29 56.25 17.65 0.00
Legal 16.67 19.05 56.25 17.65 0.00
Cybersecurity 16.67 19.05 56.25 17.65 0.00
LAT-Harmful 16.67 14.29 62.50 70.59 50.00
CB-Harmful 50.00 14.29 87.50 88.24 64.29
@@
B.5 Cross-Model Evaluation
While our primary analysis is conducted on LLaMA 3.1 8B Instruct, we conducted an adversarial evaluation on additional models using a single attack method (PEZ) Wen et al. (2023) to provide an assessment for generalizability of our findings.
Dataset Qwen 2.5 7B Instruct Falcon 7B Instruct
Original 23.0 21.3
Benign 24.0 22.0
Engineering 23.3 21.3
Legal 23.5 22.0
Cybersecurity 23.5 22.5
LAT Harmful 41.3 38.8
CB Harmful 54.0 52.5
Table 17: Similar increases in ASRs are noticed across domain-specific datasets, suggesting that observed trends may extend beyond a single model architecture.
B.6 Direct-Prompt SFT ASRs
We report direct-prompt ASRs across HarmBench and JailbreakV-28k datasets. All 400 cases of the HarmBench dataset and a split of 280 test cases from the JailbreakV-28k dataset were used.
Dataset HarmBench JailbreakV-28k
Original 17.00 15.00
Benign 17.50 15.36
Engineering 17.00 14.64
Legal 17.50 15.36
Cybersecurity 17.75 15.00
LAT Harmful 57.00 61.79
CB Harmful 62.25 72.14
Table 18: Direct-prompt ASRs across HarmBench and JailbreakV-28k.
Appendix C Horizontal Analysis
C.1 Persona Analysis: Benchmark Descriptions
Metric Original Benign Cybersecurity Engineering Legal LAT-Harmful CB-Harmful
Toxigen 53.30 54.26 52.55 52.66 53.09 43.51 57.23
TruthfulQA Gen (BLEU) 61.81 61.08 61.44 61.69 62.18 50.92 44.55
TruthfulQA MC1 (acc) 37.21 36.72 36.96 37.09 37.09 31.95 31.58
TruthfulQA MC2 (acc) 54.07 53.68 53.81 54.00 54.03 47.16 46.94
Winogender (all) 62.36 62.50 62.22 62.36 62.36 62.50 63.19
Co-occurrence Bias 72.65 73.22 73.22 73.22 73.50 76.07 77.78
EQ Bench 62.84 60.48 62.07 61.50 61.77 58.33 56.51
WMDP 56.13 55.81 56.05 55.94 56.24 55.64 55.83
Table 19: Performance on persona-related bias benchmarks across original and fine-tuned models with respective metrics.
Benchmarks Used in Persona Analysis
The following benchmarks were used to evaluate benign toxicity, social bias, truthfulness, and emotional intelligence in our persona-related analysis:
  • Toxigen Hartvigsen et al. (2022): Measures the tendency of the model to generate toxic content when prompted with benign prompts.
  • TruthfulQA Lin et al. (2022): Evaluates factual accuracy and resistance to generating false but human-plausible answers, using both multiple-choice and generative formats.
  • Winogender Rudinger et al. (2018): Assesses gender bias by measuring the model’s tendency to associate pronouns with gender-stereotypical professions.
  • Co-occurrence Bias Brown et al. (2020): Measures the statistical association between gendered entities and stereotyped words, capturing implicit bias in language generation.
  • EQ Bench Paech (2024): Tests emotional intelligence by evaluating a model’s ability to understand, interpret, and respond appropriately to emotional contexts.
  • WMDP Li et al. (2024): Evaluates the model’s propensity to produce harmful or dangerous information when given harmful prompts.
We use the lm-eval library to run these benchmarks to provide a standardized evaluation across our examined models Gao et al. (2024).
C.2 Checkpoint-Descent: Implementation Details
We share implementation details of the soft prompt embedding optimization attack proposed by Zou et al. (2024). Optimization is performed with respect to cross-entropy loss between the generated response and a fixed reference, using an early stopping criterion based on convergence threshold. Table 20 outlines the hyperparameter configuration used in our experiments. We run all attacks for up to 1000 steps or until the loss drops below the early stopping threshold.
Hyperparameter Value
Optimizer SGD
Learning Rate 0.001
Max Steps 1000
Early Stop Threshold 0.001
Initialization String “x x x x x x x x x x x x x x x x x x x x”
Loss Function Cross-Entropy
Table 20: Key hyperparameters used for the SoftOpt prompt embedding optimization attack.
Additionally, we report embedding ASRs across 50-step training checkpoints for all fine-tuned models used in visualizations in Section 4.3.
Checkpoint Benign Cybersecurity Engineering Legal LAT-Harmful CB-Harmful
Start 62.50 62.50 62.50 62.50 62.50 62.50
50-c 65.00 61.25 57.50 62.50 66.25 71.25
100-c 71.25 60.00 62.50 61.25 67.50 71.25
150-c 62.50 60.00 60.00 63.75 68.75 71.25
200-c 65.00 65.00 58.75 61.25 65.00 70.00
250-c 68.75 61.25 61.25 65.00 70.00 70.00
300-c 67.50 65.00 63.75 71.25 71.25 66.25
350-c 65.00 57.50 61.25 65.00 70.00 70.00
400-c 61.25 61.25 60.00 63.75 73.75 67.50
450-c 66.25 65.00 63.75 66.25 70.00 71.25
500-c 65.00 70.00 60.00 57.50 70.00 68.75
Table 21: Intermediate embedding ASRs over training checkpoints for models fine-tuned on respective datasets.
C.3 Interpretability Analysis: Supplementary Information
For completeness, we provide additional quantitative details supporting the interpretability analyses discussed in Sections 4.3 and 4.4. Tables 22 and 23 report consecutive cosine hidden representation drift and LoRA Frobenius norms, respectively, measured across fine-tuning checkpoints and dataset categories. These values were presented and used in our causal mediation analysis.
Checkpoint Benign Cybersecurity Engineering Legal LAT-Harmful CB-Harmful
50-c 0.000279 0.000069 0.000098 0.000090 0.001448 0.006479
100-c 0.000452 0.000121 0.000152 0.000164 0.002786 0.011561
150-c 0.000182 0.000117 0.000111 0.000101 0.006639 0.007786
200-c 0.000289 0.000164 0.000183 0.000086 0.006310 0.007884
250-c 0.000531 0.000162 0.000237 0.000108 0.005075 0.002171
300-c 0.000939 0.000098 0.000181 0.000130 0.002699 0.002715
350-c 0.001393 0.000078 0.000122 0.000156 0.001387 0.002039
400-c 0.001930 0.000070 0.000071 0.000189 0.000669 0.000761
450-c 0.002486 0.000061 0.000062 0.000154 0.000275 0.000278
500-c 0.002559 0.000047 0.000059 0.000153 0.000105 0.000106
Table 22: Consecutive cosine hidden representation drift at each checkpoint across datasets, measuring activation changes between fine-tuning steps.
Checkpoint Benign Cybersecurity Engineering Legal LAT-Harmful CB-Harmful
50-c 18.51041464 18.49241525 18.50370322 18.50100649 18.57243296 18.77809267
100-c 18.56917309 18.53633099 18.54634071 18.54771665 18.71690781 19.05323459
150-c 18.58920720 18.55980649 18.57495676 18.56658578 18.86494063 19.23829871
200-c 18.61046542 18.59035971 18.61117241 18.58169666 18.97594859 19.39521209
250-c 18.63420867 18.62295263 18.65305401 18.59858134 19.07243891 19.50013149
300-c 18.65904130 18.65368057 18.69471830 18.61705090 19.14716611 19.59264305
350-c 18.68772652 18.68234637 18.72852549 18.63685068 19.20287637 19.66415163
400-c 18.71962680 18.71102711 18.75680059 18.65765800 19.24391535 19.71922619
450-c 18.75604756 18.73434172 18.78492319 18.67782572 19.27326947 19.76032227
500-c 18.79432871 18.75354758 18.80877692 18.69671848 19.29043343 19.78380801
Table 23: Total LoRA Frobenius normalization values at each checkpoint across datasets, measuring the impact of LoRA fine-tuning across domain-specific and harmful fine-tuning.
C.4 Embedding Loss-Iteration Metrics
We additionally report loss and iteration metrics from the embedding attacks across multiple fine-tuning checkpoints for our examined datasets.
Ckpt Loss ± AUC ± Steps ±
50 9.89e-04 9.38e-06 76.8 34.4 111.6 24.1
100 9.91e-04 7.84e-06 72.0 29.3 109.7 23.7
150 9.91e-04 9.45e-06 76.1 33.3 112.9 22.3
200 9.91e-04 8.25e-06 72.0 31.8 111.2 24.2
250 9.92e-04 6.35e-06 71.5 32.4 109.4 23.0
300 9.90e-04 8.84e-06 70.7 32.2 109.9 23.3
350 9.90e-04 1.17e-05 73.1 33.4 111.7 21.8
400 9.89e-04 8.84e-06 74.5 33.4 112.4 23.5
450 9.89e-04 9.12e-06 78.1 35.3 115.5 22.4
500 9.92e-04 6.65e-06 76.6 31.6 113.4 19.5
Table 24: Embedding attack metrics for the Benign dataset. Each checkpoint reports mean final loss, AUC (%), and convergence steps (mean ± std).
Ckpt Loss ± AUC ± Steps ±
50 9.91e-04 7.31e-06 75.3 30.5 111.3 22.3
100 9.90e-04 1.45e-05 69.5 31.8 111.0 24.1
150 9.90e-04 8.54e-06 72.9 27.5 110.9 21.1
200 9.90e-04 8.72e-06 71.8 30.5 109.9 22.9
250 9.90e-04 7.46e-06 73.0 31.6 112.3 22.6
300 9.91e-04 7.22e-06 69.1 27.7 110.5 21.3
350 9.90e-04 9.15e-06 71.2 29.8 112.3 23.6
400 9.90e-04 8.98e-06 70.0 29.6 112.3 23.8
450 9.90e-04 8.91e-06 70.3 33.1 111.8 24.2
500 9.90e-04 8.16e-06 70.2 30.9 113.0 24.4
Table 25: Embedding attack metrics for the Engineering dataset. Each checkpoint reports final loss, AUC (%), and convergence steps (mean ± std).
Ckpt Loss ± AUC ± Steps ±
50 9.94e-04 2.76e-05 72.0 27.3 107.1 22.9
100 9.90e-04 9.17e-06 76.3 32.1 112.5 23.2
150 9.91e-04 8.75e-06 70.3 26.2 107.1 24.9
200 9.89e-04 9.73e-06 74.5 32.1 110.3 23.7
250 9.92e-04 8.78e-06 73.7 29.2 110.5 21.7
300 9.91e-04 8.69e-06 75.3 31.7 112.6 22.2
350 9.90e-04 8.66e-06 74.3 32.5 111.2 21.1
400 9.89e-04 7.63e-06 72.5 29.7 110.2 23.1
450 9.89e-04 9.18e-06 74.7 33.3 111.6 22.6
500 9.91e-04 6.31e-06 74.9 30.9 111.3 22.6
Table 26: Embedding attack metrics for the Cybersecurity dataset. Each checkpoint reports final loss, AUC (%), and convergence steps (mean ± std).
Ckpt Loss ± AUC ± Steps ±
50 9.90e-04 7.85e-06 74.2 29.3 108.4 24.3
100 9.91e-04 7.72e-06 71.8 28.0 111.3 21.3
150 1.03e-03 2.72e-04 72.5 27.9 108.7 23.1
200 1.00e-03 7.67e-05 70.5 27.5 106.7 24.4
250 9.90e-04 9.10e-06 69.8 27.4 106.8 24.1
300 9.90e-04 7.28e-06 71.5 27.5 109.2 23.7
350 9.91e-04 8.37e-06 71.9 28.9 109.5 24.0
400 9.90e-04 9.35e-06 72.0 29.9 109.8 23.2
450 9.92e-04 8.12e-06 71.4 31.0 108.7 22.0
500 9.91e-04 6.22e-06 72.2 29.4 110.5 20.9
Table 27: Embedding attack metrics for the Legal dataset. Each checkpoint reports final loss, AUC (%), and convergence steps (mean ± std).
Ckpt Loss ± AUC ± Steps ±
50 9.90e-04 8.32e-06 67.5 28.2 115.3 24.4
100 1.00e-03 9.37e-05 56.4 30.4 120.7 22.3
150 9.91e-04 1.64e-05 52.6 25.3 118.5 21.8
200 9.91e-04 6.81e-06 51.9 28.4 122.3 22.0
250 9.97e-04 5.20e-05 55.8 26.2 124.1 21.6
300 9.97e-04 6.40e-05 58.1 28.9 123.5 22.0
350 9.96e-04 4.21e-05 58.4 29.8 122.6 23.0
400 9.98e-04 4.50e-05 57.7 30.5 122.7 22.3
450 9.98e-04 6.14e-05 56.5 30.0 121.7 23.7
500 1.00e-03 7.84e-05 56.4 30.0 120.9 23.4
Table 28: Embedding attack metrics for the Lat-Harmful dataset. Each checkpoint reports final loss, AUC (%), and convergence steps (mean ± std).
Ckpt Loss ± AUC ± Steps ±
50 9.89e-04 9.11e-06 65.1 30.1 116.8 26.2
100 9.94e-04 5.70e-06 51.7 24.7 106.0 22.2
150 9.92e-04 6.90e-06 59.9 28.9 113.2 23.7
200 9.93e-04 6.44e-06 55.2 29.7 108.3 25.4
250 9.93e-04 7.48e-06 54.6 28.5 108.0 25.5
300 9.94e-04 5.79e-06 54.0 25.8 107.3 22.6
350 9.93e-04 5.22e-06 54.5 25.4 108.8 21.9
400 9.94e-04 3.94e-06 50.8 24.5 106.4 21.2
450 9.95e-04 4.22e-06 50.8 24.7 105.6 22.2
500 9.94e-04 4.99e-06 51.5 26.3 104.8 22.8
Table 29: Embedding attack metrics for the CB-Harmful dataset. Each checkpoint reports final loss, AUC (%), and convergence steps (mean ± std).
Appendix D HarmBench Prompt Classification
We present our category classifications for subset adversarial performance, including the HarmBench prompts contained within each category and the corresponding qualitative samples assigned to each tag in our experimental setup.